Use an algebraic equation to find the measures of the two angles described below. Begin by letting X represent the degree measure of the angle's complement.
The measure of an angle is 6 degrees greater than the complement.
What is the measure of the complement?

Answer by mark_henrhiz(40)

I hope you tried it first before sending it here.
let x be the angle
let 90 - x be the complement
x = 6 + (90 - x)
90 = 6 + (90 -x) + 90 - x
90 = 6 + 90 +90 - x - x
90 = 186 - 2x
90 - 186 = - 2x
- 96 = - 2x
48 = x
90 - x = 42
the measure of the complement is 42